# PSBD: Prediction Shift Uncertainty Unlocks Backdoor Detection

### v1

@@ Abstract
Deep neural networks are susceptible to backdoor attacks, where adversaries manipulate model predictions by inserting malicious samples into the training data. Currently, there is still a lack of direct filtering methods for identifying suspicious training data to unveil potential backdoor samples. In this paper, we propose a novel method, Prediction Shift Backdoor Detection (PSBD), leveraging an uncertainty-based approach requiring minimal unlabeled clean validation data. PSBD is motivated by an intriguing Prediction Shift (PS) phenomenon, where poisoned models’ predictions on clean data often shift away from true labels towards certain other labels with dropout applied during inference, while backdoor samples exhibit less PS. We hypothesize PS results from neuron bias effect, making neurons favor features of certain classes. PSBD identifies backdoor training samples by computing the Prediction Shift Uncertainty (PSU), the variance in probability values when dropout layers are toggled on and off during model inference. Extensive experiments have been conducted to verify the effectiveness and efficiency of PSBD, which achieves state-of-the-art results among mainstream detection methods. Codes are available at https://github.com/WL-619/PSBD.
@@ 1 Introduction
The proliferation of Deep Neural Networks (DNNs) has heralded a new era in artificial intelligence, driving progress across diverse sectors, including autonomous driving and healthcare personalization [1, 2]. Yet, as their application scope broadens, DNNs have become increasingly vulnerable, particularly from a security standpoint. One of the most notable threats in this arena is the rise of backdoor attacks [3, 4, 5, 6]. These attacks involve the surreptitious insertion of altered samples into training data, enabling attackers to subtly manipulate a DNN’s output, leading to incorrect predictions under certain triggers. The implications of such vulnerabilities are especially severe in contexts demanding high security or safety, as they can lead to catastrophic outcomes [7, 8].
@@
Notwithstanding an increased recognition of these risks, there are many different type of defense strategies currently, such as backdoor model reconstruction [9, 10, 11, 12], backdoor model detection [13, 14, 15, 16], and poison suppression [17]. However, the majority of these methods primarily focus on determining whether trained models contain backdoors or on mitigating the impact of such vulnerabilities. In contrast, there is a noticeable shortage of advanced and efficient approaches that can proactively identify backdoor training samples at the initial stages. Current research in identifying backdoor training data often suffers from either a low true positive rate - indicating a low detection rate of backdoor training data, or a high false positive rate - indicating a high error rate in identifying clean training data [18, 19, 20, 21], as detailed in Table 1. This issue is largely attributed to the focus of most existing research on data-level operations without utilizing the inherent properties of the models themselves.
Figure 1: An overview of the PSBD framework.
@@
To address this identified gap, we offer a new perspective - the model predictive uncertainty and propose a novel backdoor data detection approach named Prediction Shift Backdoor Detection (PSBD), which is inspired by an intriguing Prediction Shift (PS) phenomenon. The PS phenomenon is observed when the predictions made by a poisoned model on clean data tend to deviate from the correct labels, moving towards other certain labels, especially when dropout is used during inference. Conversely, the predictions on backdoor data generated by both classical and advanced attacks remain relatively stable. This observation of PS led us to hypothesize the existence of a weights-neuron bias in DNN models, which we called the “neuron bias” effect, where some certain paths in the network become predisposed towards the specific class after training, and the backdoor samples have different paths compared with clean data. Driven by these insights, Prediction Shift Uncertainty (PSU) is designed to measure the strength of PS that computes the variability in prediction confidences when a model evaluates a sample with both enabled and disabled dropout. A lower PSU value indicates a higher likelihood of the sample being malicious. By calculating PSU, we can effectively segregate backdoor data from clean data using a small set of label-free clean validation data, thus laying the groundwork for the PSBD approach. Figure 1 provides an overview of the PSBD framework. Our approach represents a significant stride in backdoor data detection. Unlike existing methods, it focuses on the inherent uncertainty within the model, analyzing how dropout influences prediction probabilities of clean and backdoor samples.
To summarize, our main contributions are four-fold:
@@
  • We present a novel insight into the vulnerability of DNNs to backdoor attacks, linking it to the model’s inherent predictive uncertainty. Our analysis delves into the impact of dropout on PS and introduces the concept of“neuron bias” within DNN models.
@@
  • We conduct extensive experiments across multiple benchmark datasets, rigorously evaluating our method under diverse attack scenarios and comparing it with a variety of defenses, thereby demonstrating its effectiveness and robustness.
@@ 2 Related Work
Backdoor Attacks
Backdoor attacks are particularly dangerous, injecting triggers into a target model that cause it to misclassify inputs containing these triggers while operating normally on unaltered samples [3, 22, 23]. Initial approaches to backdoor attacks, such as BadNets [3] and Blend attacks [23], involved embedding obvious trigger patterns like square patches into the input data. These methods evolved into more covert techniques, like clean-label attacks [24], which subtly poison samples of the target class using adversarial methods without obvious label changes, enhancing their stealthiness. Recent advancements have led to even more refined attacks, like WaNet [5], which introduces triggers that are specific to individual samples and less noticeable.
Backdoor Defenses
In response to backdoor attacks, a multitude of defense strategies has been developed by researchers. These include efforts for backdoor trigger recovery [25, 26, 27, 28, 29], which focus on identifying and reverse-engineering the attacker’s trigger, and strategies for backdoor model reconstruction [10, 11, 12], aimed at purging the backdoor model of its malicious elements. Methods for model detection [13, 14, 15, 16] are employed to ascertain whether a model has been tainted with backdoor samples. Despite the diversity and innovation in these defenses, they generally do not address the detection of backdoor data. Techniques targeting backdoor data detection primarily concentrate on altering input data [18, 19], uncovering input masks [20], or distinguishing the feature representations of backdoor and benign samples [21]. Nevertheless, these methods consider varying inputs and often experience low detection rates of backdoor training data or high error rates in identifying clean training data, as both clean and backdoor features can either remain intact or disappear when inputs are scaled. Our paper highlights the shortcomings of relying on input data variability and introduces a novel detection method that leverages model-level uncertainty, thereby surpassing the performance of methods based on input uncertainty.
@@ 4 Method
In this section, we offer a new perspective on the inherent predictive uncertainty within the model for the vulnerability of DNNs to backdoor attacks. We begin with two pilot studies to explore the predictive uncertainty of the model on the clean data and backdoor data. Then, we present our detection method, Prediction Shift Backdoor Detection (PSBD).
@@ 4.1 A Spark of Inspiration: MC-Dropout Predictive Uncertainty
We use a widespread model predictive uncertainty approximation method - Monte Carlo Dropout (MC-Dropout) [33] to explore the model predictive uncertainty of the three types of data - the clean training data, the backdoor training data, and the clean validation data. MC-Dropout opens dropout during the inference phase, allowing the model to undergo multiple forward passes. The average of the predicted confidences generated during this process is considered the final predicted confidence of the model, while the standard deviation of the highest predicted confidence serves as the predictive uncertainty for the corresponding data.
Applying MC-Dropout to measure uncertainty, the expected result is that the clean validation data exhibits the highest uncertainty because it is not seen during the model’s training process. The uncertainty of the clean training data follows that of the clean validation data, but the gap between them should be narrower than the gap between the clean training data and the backdoor training data. This is because enabling dropout slightly impairs the image features, despite the model generally fitting the training data better than the testing data. The backdoor training data uncertainty should be the lowest because backdoor data is easier for the model to learn or remember compared to clean data. If we obtain these ideal results, based on uncertainty, we can then view the backdoor training data as outliers and use some outlier detection methods to identify the backdoor training data.
Settings.
We adopt BadNets [3] and WaNet [5] as examples for our discussion, representing visible and non-visible backdoor attacks, respectively. We conduct experiments on the CIFAR-10 dataset [34] and ResNet-18 [35], trained for 100 epochs. For both attacks, we set the poisoning ratio to 10%, i.e. replaced 10% of total training data with malicious backdoor training data. We randomly select the clean validation dataset from the original CIFAR-10 test set, and its size is 5% of the total size of the training set (including both clean training data and poisoned training data). We calculate the average MC-Dropout uncertainty of three types of data with models obtained from all 100 epochs to compare the difference in their uncertainty. To retain the setting consistent with the MC-Dropout, we place the dropout layer after each convolutional layer of every basic block in ResNet. Without sacrificing generality, the target class $y_{t}$ of backdoor data is class 0 in our all examples.
Results.
(a) Benign Model
(b) BadNets
(c) WaNet
@@ Results.
Figure 2: The average MC-Dropout uncertainty of clean training data, backdoor training data, and clean validation data under benign and poisoned models.
As shown in Figure 2(a), under the benign model, the average MC-Dropout uncertainty of backdoor training data is slightly higher than that of clean training data, both lower than clean validation data. This indicates that the benign model treats backdoor data as a product of clean data with a certain perturbation, and this perturbation does not significantly impact the benign model’s decision-making behavior. From Figure 2(b) it can be observed that, across 100 epochs, the average uncertainty of backdoor training data is significantly smaller than that of clean training data and clean validation data under BadNets, and this difference tends to stabilize in the later stages of model training. The observations align with the part of our ideal condition in the sense that backdoor examples indeed have smaller uncertainty, while clean training and clean validation data exhibit larger uncertainty. However, in Figure 2(c), one can observe that the uncertainty of backdoor training data under WaNet is not necessarily smaller than clean training data and clean validation data. The clean training data uncertainty is still lower than the clean validation data uncertainty, but the backdoor training data uncertainty is very close to or even exceeds them at some epochs. This suggests that using uncertainty defined by standard deviation alone may not be sufficient to detect backdoor data in various attack scenarios. In addition, the dropout rate $p$ is a parameter that is difficult to ascertain when we have a lack of knowledge about backdoor attacks. Other poisoned model results can be found in the Appendix A.1. We attempted a combined method with SCALE-UP (SCP) [18], which achieves better results in most scenarios. However, it still fails in some specific cases. For detailed results, please refer to Appendix  A.2.
4.2 The Enlightening Eureka Moment: Prediction Shift Phenomenon
Prediction Shift.
To delve deeper into how dropout affects the predictive uncertainty of the model, we examined how enabling dropout during the model’s forward process alters the classifications and the model’s prediction confidence. We define Prediction Shift (PS) as the phenomenon where the class predicted by the model changes before and after dropout is enabled, for samples $\mathbf{x}$ in the dataset $\mathcal{D}$. The shift ratio $\sigma$ represents the frequency of PS occurring in all forward inferences with dropout activated across the entire dataset $\mathcal{D}$, i.e.,
$\begin{split}\phi_{PS}(\mathbf{x})=\mathbb{I}\left(\mathcal{Y}(\mathbf{x};{\boldsymbol{\theta}})\neq\mathcal{Y}(\mathbf{x};{\boldsymbol{\theta}^{{}^{\prime}}})\right),\ \sigma(\mathcal{D})=\frac{1}{k|\mathcal{D}|}\sum\limits_{\mathbf{x}\in\mathcal{D}}\phi_{PS}(\mathbf{x})\end{split}$ (1)
Settings.
Results.
@@
From the below row of Figure 3(a), among all the clean training data where PS occurs, approximately 60% of them shift from the ground truth to class 3 under the benign model. The x-axis represents the shifted labels of samples that experienced a prediction shift after applying dropout. The same pattern is observed in the backdoor training and clean validation data. We observe the presence of PS even in the benign model, indicating that the PS is generated by the model during training and is a universally existing internal characteristic of the DNNs. From the above row of Figure 3(a), both clean training data and backdoor training data exhibit nearly identical shift ratio trends. This result aligns with the conclusion in the Section 4.1, i.e, the benign model treats backdoor data as a product of clean data with a certain perturbation, and classifies them mainly based on the extracted natural image features rather than the perturbation.
@@
As illustrated in Figure 3(b) and  3(c), in the BadNets and WaNet scenarios, we observe that the shift ratio curve for clean data still follows an increasing trend as $p$ increases, eventually stabilizing. However, when $p$ reaches a certain special value, the $\sigma$ for backdoor data approaches 0, while the $\sigma$ for clean data reaches a relatively high value (around 0.8). The most important thing is, among the samples experiencing PS, almost all clean data shifts to the target class $y_{t}$ (class 0 in our experiments). The same phenomenon has been observed in other attack scenarios on CIFAR-10 as well. This indicates that training with backdoor samples enhances the PS phenomenon of clean data while suppresses that of backdoor data. This is likely due to significant differences in the internal behavior of the model towards clean data and backdoor data under appropriate dropout $p$. Other poisoned model’s shift ratio curve and shift intensity can be found in the Appendix B.1.
@@
In addition, we continued to observe similar patterns on the expanded and intricate Tiny ImageNet dataset. However, the shift classes of the clean training data and the clean validation data exhibit a predominant inclination towards a certain class rather than the target class. Nevertheless, there is still a certain proportion that exhibits a bias towards the target class. Despite assuming that the defender can freely choose the model architecture, we also conducted experiments using the VGG to demonstrate that our method is not dependent on any specific model architecture. For all results, please refer to the Appendix B.2 and  B.3.
Neuron Bias.
We posit that the PS phenomenon arises due to the “neuron bias” effect in the network during training with data, where neurons become predisposed towards features highly representative of the certain class implicitly. This bias intensifies as the network forms a strong association between the data features and this certain class, especially between the backdoor pattern and the target class. Once this association is established, the bias persists independently of the trigger pattern, essentially creating durable pathways to the target class. In the absence of dropout, models usually predict the actual class for clean data as the features corresponding to their ground truth classes outweigh the influence of these robust pathways. However, under dropout conditions, these distinguishing features of clean data are diminished, leaving a part of the robust pathways unchallenged and leading to predictions favoring the certain class, especially the target class when the model has good generalization ability as these pathways can be established with greater stability. Therefore, even with clean data input, the network tends to favor the target class.

### v2

@@ Abstract
Deep neural networks are susceptible to backdoor attacks, where adversaries manipulate model predictions by inserting malicious samples into the training data. Currently, there is still a significant challenge in identifying suspicious training data to unveil potential backdoor samples. In this paper, we propose a novel method, Prediction Shift Backdoor Detection (PSBD), leveraging an uncertainty-based approach requiring minimal unlabeled clean validation data. PSBD is motivated by an intriguing Prediction Shift (PS) phenomenon, where poisoned models’ predictions on clean data often shift away from true labels towards certain other labels with dropout applied during inference, while backdoor samples exhibit less PS. We hypothesize PS results from the neuron bias effect, making neurons favor features of certain classes. PSBD identifies backdoor training samples by computing the Prediction Shift Uncertainty (PSU), the variance in probability values when dropout layers are toggled on and off during model inference. Extensive experiments have been conducted to verify the effectiveness and efficiency of PSBD, which achieves state-of-the-art results among mainstream detection methods. The code is available at https://github.com/WL-619/PSBD.
@@ 1 Introduction
The proliferation of Deep Neural Networks (DNNs) has heralded a new era in artificial intelligence, driving progress across diverse sectors, including computer vision, autonomous driving and healthcare personalization [24, 36, 31, 41]. Yet, as their application scope broadens, DNNs have become increasingly vulnerable from a security standpoint. One of the most notable threats in this arena is the rise of backdoor attacks [13, 10, 32, 7]. These attacks involve the surreptitious insertion of altered samples into training data, enabling attackers to subtly manipulate a DNN’s output, leading to incorrect predictions under certain triggers. The implications of such vulnerabilities are especially severe in contexts demanding high security, as they can lead to catastrophic outcomes [1, 5].
Figure 1: A simple conceptual diagram of the Prediction Shift Backdoor Detection (PSBD) framework. The introduction of dropout during the inference stage induces a neuron bias effect in the model, causing the final feature maps of clean data and backdoor data to become highly similar, ultimately leading to the occurrence of the Prediction Shift phenomenon, which serves as a basis for detecting backdoor training data.
@@
Notwithstanding an increased recognition of these risks, there are many different type of defense strategies currently, such as backdoor model reconstruction [49, 2, 20, 33, 30], backdoor model detection [46, 38, 48, 3], and poison suppression [26]. However, the majority of these methods primarily focus on determining whether trained models contain backdoors or on mitigating the impact of such vulnerabilities. In contrast, there is a noticeable shortage of advanced and efficient approaches that can proactively identify backdoor training samples at the initial stages. Current research in identifying backdoor training data often suffers from either a low true positive rate - indicating a low detection rate of backdoor data, or a high false positive rate - indicating a high error rate in identifying clean data [14, 11, 19, 4, 34], as detailed in Table 1. This issue is largely attributed to the focus of most existing research on data-level operations without utilizing the inherent properties of the models themselves. To address this identified gap, we offer a new perspective - the model predictive uncertainty and propose a novel backdoor data detection approach named Prediction Shift Backdoor Detection (PSBD), which is inspired by an intriguing Prediction Shift (PS) phenomenon.
@@
The PS phenomenon is observed when the predictions made by a poisoned model on clean data tend to deviate from the correct labels, moving towards other certain labels, especially when dropout is used during inference. Conversely, the predictions on backdoor data generated by both classical and advanced attacks remain relatively stable. This observation of PS led us to hypothesize the existence of a weights-neuron bias in DNN models, which we called the “neuron bias” effect, where some certain paths in the network become predisposed towards the specific class after training, and the backdoor samples have different paths compared with clean data. Figure 1 provides a brief overview of the PSBD framework and the neuron bias effect. Under normal conditions, the feature maps (the final convolutional layer) of clean and backdoor data exhibit significant differences. However, after applying the dropout, the neuron bias effect is induced within the model, causing the feature maps of clean and backdoor data to become strikingly similar, ultimately resulting in the occurrence of the PS phenomenon. We also provided more detailed experimental explanations for verifying the neuron bias effect in section 4.2.
Driven by these insights, Prediction Shift Uncertainty (PSU) is designed to measure the strength of PS that computes the variability in prediction confidences when a model evaluates a sample with both enabled and disabled dropout. A lower PSU value indicates a higher likelihood of the sample being malicious. By calculating PSU, the PSBD approach can effectively segregate backdoor data from clean data using a small set of label-free clean validation data.
Our approach represents a significant stride in backdoor data detection. Unlike existing methods, it focuses on the inherent uncertainty within the model, analyzing how dropout influences prediction probabilities of clean and backdoor samples. In summary, our main contributions are four-fold:
@@
  • We present a novel insight into the vulnerability of DNNs to backdoor attacks, linking it to the model’s inherent predictive uncertainty. Our analysis delves into the impact of dropout on PS and introduces the concept of neuron bias within DNN models.
@@
  • We conduct extensive experiments across multiple benchmark datasets, rigorously evaluating our method under diverse attack scenarios and comparing it with a variety of defenses, demonstrating its effectiveness and robustness.
@@ 2 Related Work
Backdoor Attacks.
Backdoor attacks are particularly dangerous, injecting triggers into a target model that cause it to misclassify inputs containing these triggers while operating normally on unaltered samples [13, 28, 6]. Initial approaches to backdoor attacks, such as BadNets [13] and Blend attacks [6], involved embedding obvious trigger patterns like square patches into the input data. These methods evolved into more covert techniques, like clean-label attacks [44], which subtly poison samples of the target class using adversarial methods without obvious label changes, enhancing their stealthiness. Recent advancements have led to even more refined attacks, like WaNet [32], which introduces triggers that are specific to individual samples.
Backdoor Defenses.
Researchers have developed various defenses against backdoor attacks. These include efforts for backdoor trigger recovery [45, 15, 29, 47, 18], which focus on identifying and reverse-engineering the attacker’s trigger, and strategies for backdoor model reconstruction [2, 20, 33], aimed at purging the backdoor model of its malicious elements. Methods for model detection [22, 46, 38, 48] are employed to ascertain whether a model has been tainted with backdoor samples. Backdoor sample detection evaluates whether a given sample triggers backdoor behavior in a model. Spectral Signatures (SS) [43] employs deep feature statistics to differentiate between clean and backdoor samples, but its robustness weakens with varying poisoning rates [16]. STRIP [11] blends potentially backdoored samples with a small subset of clean samples and then using the entropy of the predictions for detection. Scale-up (SCP) [14] identifies and filters malicious testing samples by analyzing their prediction consistency during pixel-wise amplification. These methods primarily concentrate on altering input data, uncovering input masks, or distinguishing the feature representations of backdoor and benign samples.
Nevertheless, these methods consider varying inputs and often experience low detection rates of backdoor training data or high error rates in identifying clean training data, as both clean and backdoor features can either remain intact or disappear when inputs are scaled. Our paper highlights the shortcomings of relying on input data variability and introduces a novel detection method that leverages model-level uncertainty, thereby surpassing the performance of methods based on input uncertainty.
@@ 4 Method
In this section, we offer a new perspective on the inherent predictive uncertainty within the model for the vulnerability of DNNs to backdoor attacks. We begin with two pilot studies to explore the predictive uncertainty of the model on the clean data and backdoor data. Then, we present our Prediction Shift Backdoor Detection (PSBD) method.
@@ 4.1 A Spark of Inspiration: MC-Dropout Predictive Uncertainty
(a) BadNets
(b) WaNet
@@
Figure 2: The average MC-Dropout uncertainty of clean training data, backdoor training data, and clean validation data under poisoned models.
We use a widespread model predictive uncertainty approximation method - Monte Carlo Dropout (MC-Dropout) [9] to explore the model predictive uncertainty of the three types of data - the clean training data, the backdoor training data, and the clean validation data. MC-Dropout activates dropout during inference, allowing multiple forward passes. The model’s final predicted confidence is the average of these passes, while the standard deviation of the highest confidence class indicates predictive uncertainty.
@@
When applying MC-Dropout, clean validation data should show the highest uncertainty, as it’s unseen during training. Clean training data should follow closely, with a smaller gap between them than between clean and backdoor training data. Backdoor data should have the lowest uncertainty, as it’s easier for the model to learn. If these patterns hold, backdoor training data can be treated as outliers, enabling their detection using outlier methods.
Settings.
We adopt BadNets and WaNet as examples for our discussion. We conduct experiments on the CIFAR-10 dataset [23] and ResNet-18 [17], trained for 100 epochs. For both attacks, we set the poisoning ratio to 10%, i.e. replaced 10% of total training data with malicious backdoor training data. Without sacrificing generality, the target class $y_{t}$ of backdoor data is class 0 in our all examples. We randomly select the clean validation dataset from the original CIFAR-10 test set, and its size is 5% of the total size of the training set. We calculate the average MC-Dropout uncertainty of three types of data with models obtained from all 100 epochs to compare the difference in their uncertainty.
Results.
Figure 2(a) shows that, over 100 epochs, the average uncertainty of backdoor training data under BadNets is significantly lower than that of clean training and validation data, with this difference stabilizing in later training stages. This aligns with our expectation that backdoor examples have smaller uncertainty. However, in Figure 2(b), the uncertainty of backdoor data under WaNet sometimes matches or exceeds that of clean data. Additional results are available in Appendix A.1. We also tested a variant of the MC-Dropout method, which showed some improvement in detection but still failed in certain cases (details in Appendix A.2). These findings suggest that using uncertainty based on standard deviation may be insufficient for detecting backdoor data across different attack scenarios. Additionally, determining the appropriate dropout rate $p$ is challenging without detailed knowledge of backdoor attacks.
4.2 The Enlightening Eureka Moment: Prediction Shift Phenomenon
Prediction Shift.
To delve deeper into how dropout affects the predictive uncertainty of the model, we examined how enabling dropout during the model’s forward process alters the model’s classifications and prediction confidence. We define Prediction Shift (PS) as the phenomenon where the class predicted by the model changes before and after dropout is enabled, for samples $\mathbf{x}$ in the dataset $\mathcal{D}$. The shift ratio $\sigma$ represents the frequency of PS occurring in all forward inferences with dropout activated across the dataset $\mathcal{D}$, i.e.,
$\begin{split}\phi_{PS}(\mathbf{x})&=\mathbb{I}\left(\mathcal{Y}(\mathbf{x};{\boldsymbol{\theta}})\neq\mathcal{Y}(\mathbf{x};{\boldsymbol{\theta}^{{}^{\prime}}})\right),\\ \sigma(\mathcal{D})&=\frac{1}{k|\mathcal{D}|}\sum\limits_{\mathbf{x}\in\mathcal{D}}\phi_{PS}(\mathbf{x})\end{split}$ (1)
Settings.
(a) Clean image w/o dropout
(b) Backdoor image w/o dropout
(c) Clean image w/ dropout
(d) Backdoor image w/ dropout
Figure 4: The first 64 feature maps out of the 512 extracted by the top layer of the model. The red boxes represent the feature map values are non-zero and the difference between each activation value in the clean and backdoor feature maps is no greater than 1. The features of clean and backdoor image become almost identical with dropout, verifying the existence of neuron bias effect.
Results.
@@
In the bottom row of Figure 3(a), about 60% of clean training data that experience Prediction Shift (PS) under the benign model shift to class 3. The x-axis shows the shifted labels, while the y-axis represents the intensity of the shift—the proportion of times a sample was predicted to a particular class during PS. This pattern is also seen in backdoor training and clean validation data, suggesting that PS is a universal characteristic of DNNs. In the top row of Figure 3(a), both clean and backdoor training data exhibit similar shift ratio trends, supporting the conclusion from Section 4.1 that the benign model treats backdoor data as perturbed clean data, classifying them mainly based on natural image features.
@@
As illustrated in Figure 3(b) and 3(c), in the BadNets and WaNet scenarios, we observe that the shift ratio curve for clean data still follows an increasing trend as $p$ increases, eventually stabilizing. However, when $p$ reaches a certain special value, the $\sigma$ for backdoor data approaches 0, while the $\sigma$ for clean data reaches a relatively high value (around 0.8). The most important thing is, among the samples experiencing PS, almost all clean data shifts to the target class $y_{t}$ (class 0 in our experiments). The same phenomenon has been observed in other attack scenarios on CIFAR-10 as well. This indicates that training with backdoor samples enhances the PS phenomenon of clean data while suppresses that of backdoor data. This is likely due to significant differences in the internal behavior of the model towards clean data and backdoor data under appropriate dropout $p$. Other poisoned model’s results can be found in the Appendix B.1.
@@
In addition, we continued to observe similar patterns on the expanded and intricate Tiny ImageNet dataset. However, the shift classes of the clean training data and the clean validation data exhibit a predominant inclination towards a certain class rather than the target class. Nevertheless, there is still a certain proportion that exhibits a bias towards the target class. Despite assuming that the defender can freely choose the model architecture, we also conducted experiments using the VGG[39] to demonstrate that our method is not dependent on any specific model architecture. For all results, please refer to the Appendix B.2 and B.3.
“Neuron Bias” - An Explanation to Prediction Shift.
We posit that the PS phenomenon arises from the neuron bias effect in the network during training, where neurons become predisposed to features highly representative of certain classes. This bias intensifies as the network establishes strong associations between specific data features and particular classes, especially in the case of backdoor features linked to the target class. In the absence of dropout, backdoored models typically predict the correct class for clean data, as they possess sufficient features to make accurate predictions. However, under dropout conditions, many key distinguishing features in clean data are discarded. Consequently, the model relies more heavily on the neuron bias established during training, leading it to classify clean data to the label associated with this bias. In contrast, the model learns backdoor data patterns more effectively and rapidly, resulting in a more stable and pronounced neuron bias. This enhanced bias allows the model to correctly classify backdoor data even when some features are omitted due to dropout.
To validate our hypothesis, we analyzed the features extracted by the BadNets model from both clean image and its corresponding backdoor version, comparing the results with and without the application of dropout. We presented the first 64 feature maps out of the 512 extracted by the top layer of the model. As illustrated in Figure 4(a) and 4(b), without the dropout, the features of clean and backdoor version exhibit minimal similarity, which partly explains the model’s distinct behavior towards these two types of data. However, under an appropriate dropout rate, Figure 4(c) and 4(d) clearly shows that the features of clean and backdoor version become almost identical with dropout. The red boxes in the figure highlight regions where the feature map values are non-zero and the difference between each activation value in the corresponding feature maps is no greater than 1. This finding successfully confirms the validity of our neuron bias effect hypothesis. Detailed results are available in Appendix C.